# Towards Foundation Models: Evaluation of Geoscience Artificial Intelligence with Uncertainty

### v1

@@
Abstract
Artificial intelligence (AI) has transformed the geoscience community with deep learning models (DLMs) that are trained to complete specific tasks within workflows. This success has led to the development of geoscience foundation models (FMs), which promise to accomplish multiple tasks within a workflow or replace the workflow altogether. However, lack of robust evaluation frameworks, even for traditional DLMs, leaves the geoscience community ill prepared for the inevitable adoption of FMs. We address this gap by designing an evaluation framework that jointly incorporates three crucial aspects to current DLMs and future FMs: performance uncertainty, learning efficiency, and overlapping training-test data splits. To target the three aspects, we meticulously construct the training, validation, and test splits using clustering methods tailored to geoscience data and enact an expansive training design to segregate performance uncertainty arising from stochastic training processes and random data sampling. The framework’s ability to guard against misleading declarations of model superiority is demonstrated through evaluation of PhaseNet, a popular seismic phase picking DLM, under 3 training approaches. Furthermore, we show how the performance gains due to overlapping training-test data can lead to biased FM evaluation. Our framework helps practitioners choose the best model for their problem and set performance expectations by explicitly analyzing model performance at varying budgets of training data.
@@ 1 Introduction
As geoscience has breached the big data world within the last 20 years [3], the automation of geophysical tasks [^1] In the AI community, a task is a single defined goal that a model should accomplish, such as classifying an image or acting as a chatbot. through artificial intelligence (AI) via deep learning models (DLMs) is booming [44, 23, 27, 38, 26]. Foundation models (FMs), a subset of AI models trained on vast amounts of data in a self-supervised fashion that can be adapted for various downstream tasks using relatively smaller labeled datasets [4], are extending DLMs into new horizons and are actively being developed for geoscience [37]. The technological leap from traditional geoscience DLMs to FMs is inevitable, yet we still lack robust evaluation methods for even the simpler DLMs. As a result, we are unprepared to fairly compare and evaluate geoscience FMs. Without a robust evaluation approach, we are unable to determine whether new technological developments are truly state of the art, we lack diagnostics that assist developers in pinpointing areas for model improvement, and we fail to convey the scenarios under which a model can be trusted. In this research, we address this gap by identifying three fundamental evaluation aspects for AI that are directly applicable to current DLMs and extendable to future FMs, uniting them under a single evaluation framework.
@@
The first and perhaps most important evaluation aspect is performance uncertainty. Two identical AI model architectures with identical training approaches will exhibit different performances as a result of different initializations and random batching during training [6, 35]. This is often referred to as aleatoric uncertainty. Here, we refer to it as training uncertainty or variability. Furthermore, the data sample and how data is split data into training, validation, and test partitions have drastic random effects on performance benchmarking [6]. This aspect is often referred to as epistemic uncertainty. Here, we refer to it as data uncertainty or variability. Determining algorithmic superiority in light of both sources of variation is a known issue [9, 16, 10, 40] and, to our knowledge, has not been explored for the geoscience AI community.
The second vital, yet often unexplored aspect helpful to researchers, and critical for FM comparison, is evaluating the model’s learning efficiency. Here, efficiency is defined as the relationship between the gain in performance to the additional training data required (see also [18]). Mapping this relationship helps decision makers compute the cost to benefit ratio of acquiring more data to meet a problem’s performance requirements, especially in data scarce scenarios [1]. It is particularly important to FMs because they are fine-tuned using smaller, targeted datasets for the task at hand, and the FM that best adapts may be preferred. This aspect may appear easy to measure by training two models on varying budgets of training data and comparing their performances. However, evaluating learning efficiency for geoscience domains requires thoughtful construction of these budgets, as simple random samples from geoscience datasets are often highly correlated due to geography, a limited number of measurement sites, and repeated measurements of a single source. This concept, known as data leakage, has been previously explored in the biological community [12, 33]. Removing or reducing data leakage to the extent possible when creating varying budgets of training data is crucial for unbiased estimation of learning efficiency.
The third fundamental aspect applicable to AI evaluation is the effect on model performance resulting from overlap in the training and test data. An overlap between training and test data is a clear example of data leakage that can lead to overestimation of model performance. To address this, it is common practice in the AI community to clearly partition the data into train and test sets such that evaluation on the test data reflects model’s performance on truly new (or unseen) data. However, ensuring no overlap between train and test data may be difficult for FMs, because a major component of the "foundation" in FMs is the big data used for pre-training that is often scraped from any available source. As an example, one of these sources could be the benchmark dataset that an evaluator is planning to use for testing. In comparing two FMs, one that happens to be trained on said benchmark and one that was not, we must avoid biased evaluation (see [11] for an example from the natural language processing community). Furthermore, spatial correlation in geoscience data warrants meticulous construction of training, validation, and test splits in contrast to simple random sampling for determining the effect in performance from overlap [17, 11, 12].
@@
This research fundamentally contributes an evaluation framework for geoscience models that intertwines all three aspects: performance variation from data and training, learning efficiency, and train-test overlap that is directly applicable to current DLMs and intends to prepare the geoscience community for evaluation of FMs. We utilize concepts from statistical experimental design to measure each aspect in turn. We demonstrate our framework using a seismic DLM to directly compare 3 training approaches: training from scratch and two transfer learning approaches. We measure uncertainty through deep ensembles and random data selection processes, learning efficiency through careful data splitting and varying data amounts, and train-test overlap using overlapping and non-overlapping datasets. Finally, we cast all performance results into a statistical framework for direct comparison with uncertainty.
@@
We use a conventional seismic DLM [^2] The evaluation framework was developed with specific challenges of FMs in mind. This development was forward-leaning and at the time of conducting this research, there are no geoscience FMS available for direct evaluation. Demonstration of the framework with a geoscience FM, particularly for evaluating downstream tasks, is a research path of future interest. for demonstrating the evaluation framework. A plethora of research in applying DLMs to the cardinal task of phase-picking provides a strong basis for our demonstration. Phase picking is the task of identifying and predicting the time that event signatures arrive, typically the primary and/or secondary phases (P-arrival and S-arrival, respectively), at a seismic station. Phase picking is a vital task for detecting, locating, and characterizing seismic events such as earthquakes or explosions. It can be a cumbersome task, motivating researchers to actively constructing better DL phase picking models (referred to as pickers here-on) [7, 28, 46, 5, 47, 48, 39, 31] and others are using these pickers to expedite workflows, including catalog creation [8, 32, 43, 2]. Of course, the best performing and most robust picker is preferred, but determining which picker is best for a given context remains a challenge that our framework is positioned to address.
@@
The paper is formulated as follows. We first introduce the data sets and how the data splits are constructed to incorporate the three fundamental evaluation aspects. Then, we briefly introduce the model architecture and training scheme, including curating the inputs and targets, and provide a detailed review of the 3 training approaches and 5 models used for comparison. This is followed by the evaluation approach including the statistical framework that distills the different model performances for digestible comparison. Finally, we present the results in terms of performance and uncertainty and finish with concluding remarks and future work.
@@ 2 Data
Figure 1: An example waveform from the curated INSTANCE dataset showing the 3 component response with estimated P-arrival (labeled P-pick is the vertical blue line) and estimated S-arrival (labeled S-pick is the vertical red line) times.
We use two publicly available, labeled, non-continuous waveform datasets: the Stanford Earthquake Dataset (STEAD; [29]) and the Italian seismic dataset for machine learning (INSTANCE; [25]). STEAD is used for pre-training and INSTANCE for subsequent training and testing. STEAD contains $\sim 235k$ noise waveforms and $\sim 1.03M$ source waveforms representing $\sim 441k$ unique sources. For the INSTANCE dataset, we use the version measured in counts and only the earthquake waveforms measured in meters per second (mps). The resulting INSTANCE set includes $\sim 132k$ noise waveforms and $\sim 824k$ source waveforms arising from $\sim 54k$ unique sources. An example waveform from INSTANCE is shown in Figure 1 along with the labeled P-arrival and S-arrival. We access both datasets through the publicly available Python seismic deep learning package, SeisBench [42], which also provides a more detailed description of each dataset.
To measure the effect of overlap between training and test data on performance, we utilize the geographical overlap between STEAD (a global dataset) and INSTANCE (Italy only dataset). This results in a 3rd dataset which is a subset of STEAD we will refer to as STEAD Masked. STEAD Masked uses all of STEAD except the waveforms whose source or station locations overlap with INSTANCE’s area. We define INSTANCE’s area as the rectangle whose 4 corners are the pairings of the maximum and minimum source or station latitude/longitude with a 5 degree buffer in all directions. This rectangular mask reduces STEAD by 1.72% of the noise waveforms and 2.82% of the earthquake waveforms and reduces the number of unique sources by 0.97%.
@@ 2.1 Training, Validation, and Test Splits
Default splits provided in SeisBench for STEAD are used to pre-train our models. The STEAD masking process removes waveforms from each split at an approximately equal rate of 4%, so no bias is induced. A more meticulous process is implemented for INSTANCE data since it is used for both training and testing models. In the case of INSTANCE data, simply dividing waveforms or sources randomly into training, validation, and test partitions is inappropriate because data from nearby sources or stations share similar sources, geographies, and processes, and hence, are correlated. Therefore, if we randomly sample the data on a waveform or source basis, we inadvertently add information about the other waveforms or sources within the dataset that were not selected. This results in data leakage, a challenge described in the introduction that impacts evaluation of learning efficiency as well as training and test overlap. To address this, we divide the INSTANCE dataset into spatial clusters, a simple and strong indicator of signal commonality. We have the option to cluster earthquake waveforms based on source or station locations and we choose to cluster based on source locations for three reasons:
  1. 1.
    sources are most often detected by sensor stations nearest to the source,
  2. 2.
    waveforms sharing the same source are assumed to have more in common than waveforms sharing the same sensor,
  3. 3.
    and the sources tend to group themselves based on geographic contours across Italy, thereby making a clustering approach natural.
We use a k-means algorithm from the Python package, Scikit-learn [34], to cluster earthquake sources based on their latitude and longitude. We chose $k=20$ clusters because it lead to clusters large enough to feasibly represent the ground characteristics of their geographic region, yet small enough to explore model performance in limited data settings. To assign the noise waveforms, which have no corresponding source locations, to the established clusters, we use the trained k-means model to predict its associated cluster based on its station location. The clusters and number of sources, stations and waveforms are shown in Figure 2. Table 1 summarizes the amount of data in the training, validation, and test splits and the construction of the splits are described next.
@@
To construct the INSTANCE splits, we first segregate the clusters into two groups: a training/validation pool and a test pool. We select the 12 central clusters to be the pool from which the training/validation data will be constructed, and 8 clusters - the 4 northern most and 4 southern most - to be the test clusters (see Figure 2). Choosing the test clusters from northern and southern regions allows us to explore two different concepts. First, they help to evaluate model performance on semi-out-of-distribution data: we expect the waveforms to share similar signatures as a whole because they are regionally constrained, but they should be slightly different due to more localized geological distinctions between the areas. Second, these splits represent a realistic situation where we may know how a model performs in a certain region and we want to measure its consistency for neighboring regions.
To mitigate regional bias in the the INSTANCE test split, we balance the data from north and south regions. The northern region is smaller comprising 2,269 sources, and therefore we randomly sample 2,269 sources from the southern region. For each source selected in the test split, we include all associated waveforms, resulting in a test set totaling 49,516 earthquake waveforms: 23,086 from the north and 26,430 from the south. To maintain the overall ratio of 11.4% noise to earthquake waveform ratio in INSTANCE, we randomly select 2,822 noise waveforms from each of the south and north regions.
@@
We aim for an 80-20 ratio for the training to validation splits and we desire an equal number of sources from each cluster to reduce bias. Since the smallest of the 12 central clusters has 795 sources, we randomly sample 159 sources from each cluster for the validation set. This totals to 1,908 sources encompassing 27,169 earthquake waveforms. Maintaining a $\sim 11.4\%$ earthquake to noise ratio, we then randomly sample 276 noise waveforms per cluster totaling 3312 noise waveforms for the validation split.
@@
3 Methods - Model Architecture, 3 Training Approaches, and 5 Models
Table 2: The details of the 5 models are given in each row including their names, if the model is initialized randomly or using weights from an upstream model trained on either STEAD or STEAD Masked, and whether weights were frozen.
@@ 3 Methods - Model Architecture, 3 Training Approaches, and 5 Models
TL Frozen Masked STEAD Masked Half Frozen
@@
We develop a unifying framework for evaluating AI models that considers uncertainty, learning efficiency, and training-test overlap to guide model selection. We demonstrate our framework by comparing 5 models as outlined in Table 2 to compare the effects of 3 model training approaches: training from scratch (the standard DL approach) and two transfer learning (TL) approaches with varying number of free parameters. However, the framework presented is sufficiently broad where one could explore the effects of different hyperparameters, training approaches, or new architectures altogether. We first review the model architecture and other commonalities among the models, then we review the uniqueness of each model and how the training design allows for estimation of data and training uncertainty.
@@
All models use the PhaseNet [47] deep learning architecture which has been extensively utilized for phase picking [20, 5, 30] and features a U-Net type architecture [36]. PhaseNet intakes 3 component waveforms of length 3,001 and is trained to output probabilities of P-arrivals, S-arrivals, and noise for each sample (time point in the waveform). Since our goal is to demonstrate the evaluation framework, we focus only on the P-arrival predictions and discard the S-arrival and noise predictions for all subsequent analysis.
All models use the PhaseNet [47] deep learning architecture which has been extensively utilized for phase picking [20, 5, 30] and features a U-Net type architecture [36]. PhaseNet intakes 3 component waveforms of length 3,001 and is trained to output probabilities of P-arrivals, S-arrivals, and noise for each sample (time point in the waveform). Since our goal is to demonstrate the evaluation framework, we focus only on the P-arrival predictions and discard the S-arrival and noise predictions for all subsequent analysis.
@@
To demonstrate our evaluation framework, we consider 3 training approaches: the standard approach for deep learning and two transfer learning approaches. The standard approach entails randomly initializing model weights and pulling the training, validation, and test splits all from the dataset of interest, in this case INSTANCE. Thus, the standard model represents this approach.
@@
The two transfer learning approaches are thought to be useful when a small amount of labeled data is available from the region of interest, but a large amount of labeled training data can be obtained from other regions. As such, one first trains a model with data from the other regions (we call an upstream model), then switches to data from the region of interest to continue learning, thereby transferring the accumulated knowledge to the new domain. In our case, we first train PhaseNet using STEAD or STEAD Masked and then further train using INSTANCE. We explore two options: 1) allow all weights to change during INSTANCE training phase and 2) fix half of the weights (the input and down sampling layers). The TL models that have fully free weights are denoted as Free, and those with fixed weights are denoted as Frozen.
To explore the effect of training-test overlap of interest to FM evaluation, we train two models for each of the two TL approaches by changing whether the model was initialized from the model trained on STEAD or STEAD Masked. Those that are seeded by STEAD Masked have the Masked label. This results in 4 TL models named TL Free, TL Free Masked, TL Frozen, and TL Frozen Masked. All 5 models (the Standard model and 4 TL models) are validated and tested using INSTANCE data.
@@
To evaluate learning efficiency, we train each model using 1, 3, 6, 9, and 12 clusters from the INSTANCE training clusters. To measure data uncertainty at a given quantity of clusters, we randomly select the cluster set 12 times, training the model each time. Then, to separate the data uncertainty from training uncertainty, we use deep ensembles [21, 14, 19]. That is, for each of the 12 cluster sets, we retrain the model 4 different times under different initializations (holding the training data sample fixed) to obtain 4 model instances for evaluation. To obtain different initializations for the Standard model, we simply randomly reset the weights. For the TL models which are seeded by weights from an upstream model, we first train 48 upstream models for both STEAD and STEAD Masked starting from random initializations. Then, we initialize the TL models by pulling the set of weights from one randomly selected upstream model.
We use ensembles to obtain training uncertainty for a few reasons. Fundamentally, deep ensembles are easy to understand and implement and they only require extending training time. This is in contrast to other uncertainty methods which generally attempt to approximate the posterior distribution of the prediction. These methods include Bayesian neural networks [22] and approximations thereof including dropout [13], Stochastic Weight Averaging - Gaussian (SWAG) [24] and many others [15]. However, these approaches enforce distributional assumptions of the model weights, make strong/arbitrary decisions (e.g, dropout), rarely consider different basins of attraction, and are generally more difficult to interpret. Some methods, however, such as MultiSWAG [41] (see [2] for a seismic based implementation), offer a strong balance between the complex Bayesian assumptions and the simpler empirical ensembles and merit consideration for future work.
@@
In our deep ensemble implementation, we loosely follow the snapshot ensemble approach of [19] which boasts computational advantages in finding different parameter states that minimize loss. We implement a cosine annealing learning schedule with 4 restarts with an initial learning rate of $0.01$ and final rate of 00. These learning parameters exhibited the best performance in preliminary training experiments especially for models with the smallest quantities of clusters (one cluster). However, simple rate resets routinely failed to push models with small training sets into new basins of attraction, so we altered the approach towards the classical deep ensemble method [21] by fully reprogramming the model weights at each scheduled rate reset rather than relying on rate resets alone. Depending on the model being trained, this reprogramming may be a complete randomized reset (for the OOD and Standard models) or replacing weights by randomly pulling from the 48 upstream STEAD (Masked) models (for the TL models).
@@
For ease of discussion, we define a single "model instance" as the combination of model, quantity of clusters, cluster set (random selection of those clusters), and initialization (ensemble). For example, a single model instance of the Standard model could be trained using 3 clusters randomly chosen - in reference to Figure 2 - to be clusters red, blue and light-green, and the third initialization. With 4 TL models and 1 Standard model each trained using 5 quantities of clusters over 12 random cluster sets with 4 initializations, we obtain 1,200 model instances.

### v2

# Evaluation of Seismic Artificial Intelligence with Uncertainty

@@
Evaluation of Seismic Artificial Intelligence with Uncertainty
Samuel Myren
Los Alamos National Laboratory | Virginia Tech
Los Alamos, NM | Blacksburg, VA
myrenst@lanl.gov
[…]
Los Alamos National Laboratory
Los Alamos, NM
[…]
Los Alamos National Laboratory
Los Alamos, NM
[…]
Los Alamos National Laboratory
Los Alamos, NM
[…]
Virginia Tech
Blacksburg, VA
[…]
Los Alamos National Laboratory
Los Alamos, NM
1 Abstract
Artificial intelligence has transformed the seismic community with deep learning models (DLMs) that are trained to complete specific tasks within workflows. However, there is still lack of robust evaluation frameworks for evaluating and comparing DLMs. We address this gap by designing an evaluation framework that jointly incorporates two crucial aspects: performance uncertainty and learning efficiency. To target these aspects, we meticulously construct the training, validation, and test splits using a clustering method tailored to seismic data and enact an expansive training design to segregate performance uncertainty arising from stochastic training processes and random data sampling. The framework’s ability to guard against misleading declarations of model superiority is demonstrated through evaluation of PhaseNet [1], a popular seismic phase picking DLM, under 3 training approaches. Our framework helps practitioners choose the best model for their problem and set performance expectations by explicitly analyzing model performance with uncertainty at varying budgets of training data.
Keywords: uncertainty quantification, model benchmark, deep learning, learning efficiency, phase pick, signal processing
@@ 2 Introduction
As seismology has breached the big data world within the last 20 years [2], the automation of seismic tasks through artificial intelligence (AI) via deep learning models (DLMs) is booming [3, 4, 5]. However, different published models are often trained and evaluated using datasets that differ on important characteristics, making it hard to compare them directly. While growing availability of open-source benchmark data sets and software packages, such as SeisBench [6], has enabled researchers to compare popular models on the same datasets [7], we still lack robust evaluation methods to fairly evaluate and compare DLMs. Without a robust evaluation approach, we are unable to determine whether new technological developments are truly state of the art, we lack diagnostics that assist developers in pinpointing areas for model improvement, and we fail to convey the scenarios under which a model can be trusted. In this research, we address this gap by identifying two fundamental evaluation aspects for DLMs and uniting them under a single evaluation framework.
@@
The first and perhaps most important evaluation aspect is performance uncertainty in terms of benchmarking. Two identical DLM architectures with identical training approaches will realize different performance metrics when tested on a benchmark dataset as a result of different initializations and random batching during training [8, 9]. We refer to this as training uncertainty. Furthermore, the training data sample has drastic random effects on performance metrics [8]. We refer to this as data uncertainty. Determining algorithmic superiority in light of both sources of uncertainty is a known issue [10, 11, 12, 13] and, to our knowledge, has not been explored for the seismic AI community. We note that our aim is not to quantify prediction uncertainty, which entails relating model variability in predictions to the prediction errors quantitatively, typically using statistical confidence/credible intervals.
The second vital, yet often unexplored, evaluation aspect is the model’s learning efficiency in terms of data utilization. Adding redundant data (too many similar examples) does not necessarily add new information, making model training inefficient. So, here, learning efficiency is defined as the relationship between the gain in performance as a result of the additional training data and data diversity required (see also [14]). Understanding a model’s learning efficiency helps decision makers compute the cost to benefit ratio of acquiring additional and more diverse data to meet a problem’s performance requirements, especially in data scarce scenarios [15]. It is particularly important in the context of transfer learning (TL) where models trained in one context are fine-tuned with smaller, targeted datasets for another context [16], and the models that best adapt are preferred.
Measuring learning efficiency while accounting for uncertainty may appear to be simple at first. One might train two models on different budgets of randomly sampled training data, compare their performances, and repeat under re-sampled data to account for data uncertainty. However, randomly sampling data does not take correlations between seismic data into account. Waveforms may share similar signal properties if they originate from similar sources, come from nearby stations, or follow similar paths. This is a form of data leakage [17], where one dataset contains overlapping information from another dataset. Data leakage can influence all aspects of model evaluation. For example, when identical or nearly identical data points exist in both training and test datasets, model performance can be overestimated. Failing to account for data leakage when choosing training budgets can lead to inaccurate measurements of learning efficiency. This is because as more data is added, there is no control over how much new information (or diversity) is being added. That is, without controlling for data leakage in developing training budgets, one may fail to systematically expose the models to more data diversity. Finally, failure to account for it can induce unrealistic representation of uncertainty arising from data sampling, potentially underestimating its impact by downplaying bias in the sampling procedure. In this work, we mitigate data leakage to more directly measure learning efficiency and better represent realistic data uncertainty by using stratified sampling to construct the training data budgets.
@@
This research develops an evaluation framework for seismic DLMs that considers two aspects: performance uncertainty arising from data and training, and learning efficiency. We utilize concepts from statistical experimental design to measure each aspect in turn. We demonstrate our framework using a seismic DLM to directly compare 3 training approaches: training from scratch and two TL approaches. We measure uncertainty through deep ensembles and random data selection processes and learning efficiency through careful data splitting and varying data amounts. Finally, we cast all performance results into a statistical framework for direct comparison with uncertainty.
@@
We use a popular seismic DLM for demonstrating the evaluation framework. A plethora of research in applying DLMs to the cardinal task of phase-picking provides a strong basis for our demonstration. Phase picking is the task of identifying the time that event related signatures arrive, typically the primary and/or secondary phases (P-arrival and S-arrival, respectively), at a seismic station. Phase picking is a vital task for detecting, locating, and characterizing seismic events such as earthquakes or explosions. It can be a cumbersome task, motivating researchers to actively construct better DL phase picking models (referred to as pickers hereafter) [18, 19, 20, 1, 21, 22, 23, 16, 24] and others are using these pickers to expedite workflows and create catalogs [25, 26, 27, 28]. Of course the best performing and most robust picker is preferred, but determining which picker is best for a given context remains a challenge that our framework is positioned to address.
@@
The paper is formulated as follows. We first introduce the data sets and how the data splits are constructed to incorporate the two fundamental evaluation aspects. Then we briefly introduce the model architecture and training scheme, including curating the inputs and targets, and provide a detailed review of the 3 training approaches used for comparison. This is followed by the evaluation approach including the statistical framework that distills the different model performances for digestible comparison with uncertainty. Finally, we present the results in terms of performance and uncertainty and finish with concluding remarks and future work.
@@ 3 Data
Figure 1: An example waveform from the INSTANCE dataset showing the 3-component response in counts (Z - top, N - middle, E - bottom) with the labeled P-arrival (vertical blue line) and labeled S-arrival (vertical red line).
We use two publicly available, labeled, non-continuous waveform datasets: the Stanford Earthquake Dataset (STEAD; [29]) and the Italian seismic dataset for machine learning (INSTANCE; [30]), both accessed through the Python seismic deep learning package, SeisBench. STEAD is used for pre-training and INSTANCE for subsequent fine-tuning and testing. STEAD contains $\sim 235k$ noise waveforms and $\sim 1.03M$ earthquake waveforms representing $\sim 441k$ unique sources. For the INSTANCE dataset, we use the version measured in counts (InstanceCountsCombined in SeisBench documentation). In addition, we use the instrument response deconvolved traces in velocity without redundant measurements in acceleration. The resulting INSTANCE set includes $\sim 132k$ noise waveforms and $\sim 824k$ earthquake waveforms arising from $\sim 54k$ unique sources. An example waveform from INSTANCE is shown in Figure 1 along with the labeled P-arrival and S-arrival.
@@ 3.1 Training, Validation, and Test Splits
We use the default splits provided in SeisBench for STEAD for pre-training the models, which we then fine-tune using INSTANCE data. We implement a more meticulous splitting process that is meant to reduce data leakage for INSTANCE, since we use it to measure learning efficiency with uncertainty. We construct the splits using stratified sampling on source locations. We chose location as a stratifying feature because it was an easy-to-implement first order way to segregate the data such that it captures some diversity and represents how seismic datasets are often constrained to specific geographic regions, thereby roughly mimicking a realistic data sampling procedure. For example, many of the publicly available benchmark datasets in SeisBench are sourced from specific geographical regions. In practice, one might want to consider how model performance improves with the inclusion of data from new regions, or how a model might perform when predicting activity in a new region with limited data.
To accomplish our stratification, we cluster the INSTANCE sources spatially based on their latitude and longitude using k-means algorithm from the Python package, Scikit-learn [31]. K-means is an iterative, centroid based clustering algorithm. Initially all cluster centroids are selected at random and all data points are assigned to a cluster closest to them. Next, in each iteration, a new centroid is estimated for each cluster based on the mean of all points assigned to this cluster and then these new centroids are used to assign clusters for each point (i.e., based on the nearest cluster centroid). This process continues until convergence. We choose 20 clusters because it provided clusters large enough to feasibly represent the region’s activity, yet small enough to explore model performance in limited data settings. To assign the noise waveforms, which have no corresponding source locations, to the established clusters, we use the trained k-means model to predict its associated cluster based on its station location. The clusters and number of sources, stations and waveforms are shown in Figure 2. Table 1 summarizes the amount of data in the training, validation, and test splits and the construction of the splits are described next.
@@
To construct the INSTANCE splits, we first segregate the clusters into two groups: a training/validation pool and a test pool. We select the 12 central clusters to be the pool from which the training/validation data will be constructed, and 8 clusters - the 4 northern most and 4 southern most - to be the test clusters (see Figure 2). Choosing the test clusters from northern and southern regions allows us to explore two different concepts. First, they help to evaluate model performance on semi-out-of-distribution data. Second, these splits represent a situation where we may know how a model performs in a certain region and we want to measure its consistency for neighboring regions.
To mitigate regional bias in the the INSTANCE test split, we balance the data from north and south regions. The northern region is smaller comprising 2,269 sources, so we randomly sample 2,269 sources from the southern region. For each source selected in the test split, we include all associated waveforms, resulting in a test set totaling 49,516 earthquake waveforms: 23,086 from the north and 26,430 from the south. To maintain the overall ratio of 11.4% noise to earthquake waveform ratio originally in INSTANCE, we randomly select 2,822 noise waveforms from each of the south and north regions.
@@
We aim for an 80-20 ratio for the training to validation splits and we desire an equal number of sources from each cluster to reduce bias. Since the smallest of the 12 central clusters has 795 sources, we randomly sample 159 sources from each cluster for the validation set. This totals to 1,908 sources encompassing 27,169 earthquake waveforms. Maintaining a $\sim 11.4\%$ earthquake to noise ratio, we then randomly sample 276 noise waveforms per cluster totaling 3,312 noise waveforms for the validation split.
@@
In our stratification, our aim was to choose a realistic feature that captures data diversity, mimics data sampling variability, and mitigates data leakage and thus reduces bias in measuring learning efficiency. We acknowledge that there are many other features that seismologists may use to stratify the waveform data, including time, network, geological faults or other geological features, source-station distance, magnitude, instrument type, earthquake type, signal to noise ratio (SNR), signal characteristics, and possibly many others. We also recognize that partitioning data on any single feature does not fully capture all forms of data diversity. However, we find that partitioning on source locations induces some segregation in these other data features. Figure 3 provides deeper insight into the how the geographical clustering impacts the distributions of selected metadata features derived from the INSTANCE dataset. More details and figures showing distributions of other data features and waveform signal properties of the clusters are provided in the appendix (Figures A1-A5). Other details regarding INSTANCE data, including geographical visualizations of source focal mechanisms and magnitudes, can be found in [30].
Figure 3: Distributions of selected data features (log epicentral distance, magnitude, log depth, SNR averaged over the 3 components, back azimuth, and source year) for each of the 12 training clusters (non-black curves corresponding to Figure 2) and the testing set comprising the 4 northern and 4 southern clusters (black curve). These features are provided as metadata in the INSTANCE dataset found in SeisBench. All curves are normalized for comparison such that they integrate to 1.
4 Methods - Model Architecture & 3 Training Approaches
Table 2: The details of the 3 models are given in each row including their names, if the model is initialized randomly or using weights from an upstream model trained on STEAD, and whether weights were frozen.
@@ 4 Methods - Model Architecture & 3 Training Approaches
We develop a unifying framework for evaluating AI models that considers performance uncertainty and learning efficiency to guide model selection. We demonstrate our framework by comparing 3 training approaches outlined in Table 2: training from scratch (the standard DL approach) and two TL approaches with varying number of free parameters. However, the framework presented is sufficiently broad where one could explore the impacts of any aspect of interest, such as different hyperparameters, training approaches, or new architectures altogether. We first review the model architecture and other commonalities among the models, then we review the uniqueness of each model and how the training design allows for estimation of data and training uncertainty.
@@
All models use the PhaseNet deep learning architecture which has been extensively utilized for phase picking [7, 32, 16] and features a U-Net type architecture [33]. PhaseNet intakes 3-component waveforms of dimension 3,001 and is trained to output probabilities of P-arrivals, S-arrivals, and noise for each time point in the waveform.
We curate the training inputs by applying a 30s window randomly around the 100 Hz sampled 3-component waveform. For earthquake waveforms, we require the window to include the labeled P-arrival. The windows are re-randomized every epoch. Then, each waveform component is independently zero-meaned and scaled to unit variance. Due to the data curation of STEAD and INSTANCE, the P-arrivals tend to appear early in the waveforms, most often within the first 15s. As such, the random windowing reduces some temporal bias, but some remains, and this is a source of bias induced by the training process. There has been some work inspecting how windowing and overlap can impact predictions [34, 26, 35], but this is not a focus of this work. Furthermore, we recognize the challenge of waveforms containing multiple unlabeled arrivals and how data augmentation can help mitigate this issue, but no other data augmentation was implemented as our focus is in the evaluation framework as opposed to progressing picker technology.
For the training targets, we follow a process similar to [1], where we construct the P and S training targets using a normal distribution centered at the labeled picks with a standard deviation of 0.1 seconds. For the noise component, we simply subtract the P and S target components from a vector of ones. Since our goal is to demonstrate the evaluation framework, we focus only on the P-arrival predictions and discard the S-arrival and noise predictions for all subsequent analysis.
@@
To demonstrate our evaluation framework, we compare 3 training approaches: the standard approach for deep learning and two TL approaches. The standard approach entails randomly initializing model weights and pulling the training, validation, and test splits all from the dataset of interest, in this case INSTANCE. Thus, the Standard model represents this approach.
@@
The two TL approaches are thought to be useful when a small amount of labeled data is available from the region of interest, but a large amount of labeled training data can be obtained from other regions. As such, one first trains a model with data from the other regions (we call an upstream model), then switches to data from the region of interest to continue learning, thereby transferring the accumulated knowledge to the new domain. In our case, we first train PhaseNet using STEAD and then fine-tune using INSTANCE. We explore two options. The first allows all weights to change during INSTANCE training phase as similarly performed in [36]. The second fixes the weights of the input and convolution layers (i.e., the left half, or encoder-side, of the network). Previous work exploring optimal layer freezing for U-Net architectures for medical image segmentation demonstrates that this choice may be non-optimal [37], but our objective is model evaluation, not finding the optimal approach for TL. The TL models that have fully free weights are denoted as Free, and those with some fixed weights are denoted as Frozen. All 3 models (the Standard model and 2 TL models) are validated and tested using INSTANCE data.
@@
To evaluate learning efficiency, we train each model using 1, 3, 6, 9, and 12 clusters from the INSTANCE training clusters. To measure data uncertainty at a given quantity of clusters, we randomly select the cluster set 12 times, training the model each time. Then, to separate the data uncertainty from training uncertainty, we use 4 deep ensembles [38, 39, 40]. That is, for each of the 12 cluster sets at a given quantity of clusters, we retrain the model 4 different times under different initializations, holding the training data sample fixed, to obtain 4 model instances for evaluation. To obtain different initializations for the Standard model, we simply randomly reset the weights. For the TL models which are seeded by weights from an upstream model, we first train 48 upstream models with STEAD starting from random initializations. Then, we initialize the TL models by pulling the set of weights from one randomly selected upstream model.
@@
In our deep ensemble implementation, we loosely follow the snapshot ensemble approach of [39] which boasts computational advantages in finding different parameter states that minimize loss. We implement a cosine annealing learning schedule with 4 restarts with an initial learning rate of $0.01$ and final rate of 00. These learning parameters exhibited the best performance in preliminary training experiments especially for models with the smallest quantities of clusters (one cluster). However, simple rate resets routinely failed to push models with small training sets into new basins of attraction, so we altered the approach towards the classical deep ensemble method [38] by fully reprogramming the model weights at each scheduled rate reset rather than relying on rate resets alone. Depending on the model being trained, this reprogramming may be a complete randomized reset (for the Standard model) or replacing weights by randomly pulling from one of the 48 upstream models (for the TL models).
@@
For ease of discussion, we define a single "model instance" as the combination of model, quantity of clusters, cluster set (random selection of those clusters), and initialization (ensemble). For example, a single model instance of the Standard model could be trained using 3 clusters randomly chosen - in reference to Figure 2 - to be clusters red, blue and light-green, and the third initialization. With the Standard model and 2 TL models each trained using 5 quantities of clusters over 12 random cluster sets with 4 initializations, we obtain 720 model instances.